OpenStudy (anonymous):

quadratic equation has one solution x=6 and x=-8 what is the equation found to be?

OpenStudy (anonymous):

x - 6

OpenStudy (anonymous):

x-8

OpenStudy (anonymous):

x+8

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

(x-6)(x+8)=0

OpenStudy (anonymous):

Add the solutions: 6+(-8) = -2 Multiply the solutions: 6*(-8) = -48 answer: \[x^2-(-2)x-48 = x^2+2x-48\]

OpenStudy (anonymous):

nice, joe

OpenStudy (anonymous):

If 'a' is the sum of the roots, and 'b' is the product of the roots, then the equation is: \[x^2-ax+b\]

OpenStudy (anonymous):

(only for quadratic equations)

OpenStudy (anonymous):

can you expand that to higher degrees, joe? :)

OpenStudy (anonymous):

if the degree's are higher, then you have to multiply the roots k at at time and sum them. For a polynomial: \[x^n+a_{n-1}x^{n-1}+a_{n-2}x^{n-2}+\ldots+a_1x+a_0\] If the roots are: \[r_1,r_2,\ldots, r_n\] Then \[a_{n-k}\] is the sum of the products of the roots taken k at a time.

OpenStudy (anonymous):

Jahtoday how is your finger ?

OpenStudy (anonymous):

lol im using 1 hand cuz im icing it

OpenStudy (anonymous):

:) nicely stated joe!

OpenStudy (anonymous):

im forgetting something....oh, (-1)^k is there as well,

OpenStudy (anonymous):

and thank you :)

OpenStudy (anonymous):

lol that is pretty nicely done